2y + 3x = 12
-4y + 5x = -2
How do I solve the above liner equations using the elimination method? Can someone kindly explain it to me thanks

Answer by mananth(16949)

2y + 3x = 12............1
-4y + 5x = -2...........2
to eliminate a variable the co-efficients of either x or y have to be same in both the equations.
If we multiply equation 1 by 2 then the co-fficient of y will be same.
multiply and add the equations
2(2y+3x)-4y+5x = 24 -2
4y+6x-4y+5x=22
11x=10
x=22/11
x=2
plug the value of x in equation 1
2y+3*2=12
2y+6=12
2y=12-6
y2y=6
y=3